### v2

# Bozen-Bolzano, IT SkillFormer

Abstract
Assessing human skill levels in complex activities is a challenging problem with applications in sports, rehabilitation, and training. In this work, we present SkillFormer, a parameter-efficient architecture for unified multi-view proficiency estimation from egocentric and exocentric videos. Building on the TimeSformer backbone, SkillFormer introduces a CrossViewFusion module that fuses view-specific features using multi-head cross-attention, learnable gating, and adaptive self-calibration. We leverage Low-Rank Adaptation to fine-tune only a small subset of parameters, significantly reducing training costs. In fact, when evaluated on the EgoExo4D dataset, SkillFormer achieves state-of-the-art accuracy in multi-view settings while demonstrating remarkable computational efficiency, using 4.5x fewer parameters and requiring 3.75x fewer training epochs than prior baselines. It excels in multiple structured tasks, confirming the value of multi-view integration for fine-grained skill assessment.
1 Introduction
Automatically estimating human expertise from video is a key challenge in computer vision, especially in domains like sports, rehabilitation, and skill training. Yet, unlike action recognition, this task requires capturing subtle differences in motion quality and precision, which are often visible only through a combination of viewpoints.
To address this, we propose SkillFormer, a lightweight and efficient architecture for multi-view skill assessment. SkillFormer processes synchronized egocentric and exocentric video streams using a TimeSformer backbone, enhanced by a novel CrossViewFusion module that leverages cross-attention and learnable gating. To enable efficient fine-tuning, it incorporates Low-Rank Adaptation (LoRA). This design achieves competitive overall accuracy while substantially improving computational efficiency through reduced trainable parameters and faster training convergence.
This work introduces the following key contributions:
  • A CrossViewFusion module that fuses multiple views via multi-head cross-attention and learnable gating, allowing the model to focus on the most informative aspects of each view.
  • A self-calibration mechanism that aligns view-specific embeddings using learnable statistics.
  • A LoRA-based fine-tuning strategy applied to the TimeSformer backbone, enabling efficient adaptation with minimal overhead.
Compared to state-of-the-art benchmarks, SkillFormer demonstrates strong overall classification accuracy in multi-view settings, which are the primary focus of our work. While our Ego-only configuration achieves comparable results to specialized single-view methods (45.9% vs 46.8%), the real strength of our approach emerges in multi-view scenarios. SkillFormer significantly improves accuracy in exocentric settings by 14% (from 40.6% to 46.3%) and in combined ego+exo setups by 16.4% (from 40.8% to 47.5%). The benefits are particularly pronounced in structured activities, where our multi-view approach achieves improvements for Basketball (77.88%), Cooking (60.53%), and Bouldering (33.52%) by effectively integrating complementary spatial and temporal features from synchronized perspectives. These substantial gains highlight the effectiveness of our integrated cross-attention fusion approach, particularly in scenarios where multiple viewpoints and fine-grained motion cues are crucial for expertise assessment.
2 Background and Related Works
Action Quality Assessment (AQA) and Proficiency Estimation focus on evaluating the quality of human actions, with applications across sports, rehabilitation, and skill training [Zhou et al.(2024)Zhou, Cai, Wang, Shum, and Liang]. Traditional approaches relied on single-modality inputs, using handcrafted features or 3D CNNs like C3D and I3D to extract spatiotemporal representations [Parmar and Morris(2019), Carreira and Zisserman(2017), Tran et al.(2015)Tran, Bourdev, Fergus, Torresani, and Paluri]. More recent methods adopt transformer-based models such as TimeSformer, Video Swin, and ActionFormer to capture long-range dependencies and fine-grained motion dynamics [Bertasius et al.(2021)Bertasius, Wang, and Torresani, Liu et al.(2021)Liu, Ning, Cao, Wei, Zhang, Lin, and Hu, Zhang et al.(2022)Zhang, Wu, and Li]. Specialized architectures like GDLT decouple grade-aware features via decoder attention [Xu et al.(2022)Xu, Zeng, and Zheng], while CoRe introduces contrastive regression for more precise skill scoring [Yu et al.(2021)Yu, Rao, Zhao, Lu, and Zhou]. Nonetheless, skill classification remains underexplored, particularly in multi-view settings.
Multi-modal AQA improves robustness by integrating complementary signals—RGB, flow, audio, depth, or pose. Skating-Mixer [Xia et al.(2022)Xia, Zhuge, Geng, Fan, Wei, He, and Zheng] and PAMFN [Zeng and Zheng(2024)] incorporate audio-visual cues via staged fusion. Others combine RGB and depth for industrial tasks [Bianchi and Lanz(2024)], or rely on skeletons to capture motion kinematics [Bianchi and Lanz(2025), Du et al.(2015)Du, Wang, and Wang, Duan et al.(2022)Duan, Zhao, Chen, Lin, and Dai]. Datasets like EgoExoLearn [Huang et al.(2024)Huang, Chen, Xu, Zhang, Yang, Pei, Zhang, Lu, Wang, Wang, and Qiao] and EgoExo4D [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray] emphasize the importance of multi-perspective inputs, but existing fusion strategies often rely on heuristic or late fusion, limiting generalization. In EgoExo4D [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray], baseline models are modality-specific—separate architectures are trained for first-person and third-person views, with predictions later combined via ensembling. These models also rely on heterogeneous pretraining datasets across modalities, making comparison and deployment under unified constraints more challenging.
Recent work on efficient adaptation, such as LoRA [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen], enables fine-tuning with fewer trainable parameters. Knowledge distillation also supports low-resource adaptation for vision tasks [Wang and Yoon(2022), Meneghetti et al.(2025)Meneghetti, Bianchi, Demo, and Rozza].
Our work advances the state-of-the-art by addressing key limitations in prior methods—such as reliance on modality-specific models, late fusion, and heavy pretraining. By introducing a unified architecture for multi-view skill classification, we simultaneously enhance classification accuracy and computational efficiency, enabling robust skill assessment under realistic resource constraints.
Figure 1: Overview of the SkillFormer architecture. Multi-view video inputs (one egocentric and up to four exocentric) are processed through a shared TimeSformer backbone fine-tuned with LoRA. Features are fused using the CrossViewFusion module and passed to a classification head.
3 Proposed Methodology
We present SkillFormer, a novel architecture for unified multi-view proficiency estimation. Our model builds upon the TimeSformer [Bertasius et al.(2021)Bertasius, Wang, and Torresani] framework, incorporating parameter-efficient fine-tuning through LoRA [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] and a specialized CrossViewFusion module for multi-view integration. Figure 1 provides an overview of our approach.
3.1 Problem Formulation
Given a set of synchronized videos from multiple viewpoints (e.g., first-person and third-person cameras) capturing a person performing an activity, our goal is to classify their expertise level into one of predefined categories (Novice, Early Expert, Intermediate Expert, or Late Expert). Formally, for each sample, we have a set of video sequences $\{V_{1},V_{2},...,V_{n}\}$ from $n$ different views, and we aim to predict a label $y\in\{0,1,2,3\}$ indicating the expertise level.
3.2 Video Feature Extraction
We use a TimeSformer backbone [Bertasius et al.(2021)Bertasius, Wang, and Torresani], pretrained on Kinetics-600 [Carreira and Zisserman(2017)], to extract spatio-temporal features by modeling spatial and temporal dependencies via divided attention. For efficient adaptation, we apply LoRA [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] to key components—including attention projections, feed-forward layers, and temporal modules—allowing us to fine-tune only a small number of additional parameters.
For multi-view inputs, we flatten the batch and view dimensions, extract features in parallel using the shared backbone, and reshape the output before fusing it with the CrossViewFusion module. This design maintains per-view identity while enabling efficient joint processing.
To support variable-length inputs while preserving alignment with the original 8-frame pretraining, we interpolate both the input frames and the model’s temporal positional embeddings. In particular, we apply linear interpolation to the learned time embeddings, enabling the model to generalize across different temporal resolutions without degrading temporal consistency.
[…]
3.3 CrossViewFusion for Unified Multi-View Integration
We introduce CrossViewFusion, a lightweight yet expressive module for fusing information from multi-view video streams. It takes as input view-specific features from egocentric and exocentric sequences and applies a sequence of operations: normalization, cross-view attention, gated transformation, and adaptive self-calibration. These steps produce a unified embedding tailored for downstream expertise classification. An overview of the module is shown in Figure 2. Below, we describe each component of the CrossViewFusion module, aligned with its actual implementation.
View-Wise Normalization: To account for statistical disparities across viewpoints, we apply layer normalization independently to each view’s features. This operation ensures consistent feature scaling and centering, which is especially critical when merging modalities with inherently different distributions, such as first-person and third-person video streams.
Multi-Head Cross-View Attention: Rather than relying on simple concatenation or averaging, we employ a multi-head cross-attention mechanism, where each view can attend to and contextualize information from all other views. This enables the model to learn inter-view dependencies and adaptively prioritize view-specific cues. For instance, in cooking scenarios, egocentric hand movements may dominate attention, whereas in dance sequences, global body posture from exocentric views may become more salient.
Gated Feature Transformation The aggregated features are passed through a feed-forward network that includes a learnable gating mechanism. This consists of a linear layer followed by a sigmoid activation that produces gating weights for each hidden dimension. The gate acts as a dynamic filter, performing an element-wise modulation of the hidden representation to selectively amplify informative features while suppressing irrelevant ones. Such modulation is crucial for expertise assessment, where fine-grained differences in motion and control—not just coarse action labels—drive prediction quality.
Adaptive Feature Self-Calibration To further align the fused representations with the target classification space, we introduce an adaptive self-calibration block that learns per-feature affine parameters ($\mu_{\text{learn}},\sigma_{\text{learn}}$). This departs from traditional fixed-statistics normalization and allows the model to reshape feature distributions in a data-driven manner, improving its ability to resolve fine-grained skill variations across diverse activities.
Classification Head Finally, the calibrated features are fed into a classifier to predict discrete expertise levels.
4 Experimental Setup
4.1 Dataset
We evaluate on the Ego-Exo4D dataset [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray], a large-scale multimodal benchmark comprising over 1,200 hours of synchronized egocentric and exocentric video from 740 participants recorded across 123 real-world environments. Egocentric videos are first-person recordings captured from the demonstrator’s viewpoint, while exocentric videos are third-person views recorded from static cameras placed in the environment. The dataset spans eight domains—cooking, health, bike repair, music, basketball, bouldering, soccer, and dance—captured in naturalistic settings to reflect authentic task execution.
For the demonstrator proficiency estimation benchmark, only six domains are used: cooking, music, basketball, bouldering, soccer, and dance. The health and bike repair domains are excluded due to highly imbalanced or incomplete label distributions.
Each sample includes one egocentric video, recorded using Project Aria glasses [Engel et al.(2023)Engel, Somasundaram, Goesele, Sun, Gamino, Turner, Talattof, Yuan, Souti, Meredith, Peng, Sweeney, Wilson, Barnes, DeTone, Caruso, Valleroy, Ginjupalli, Frost, Miller, Mueggler, Oleinik, Zhang, Somasundaram, Solaira, Lanaras, Howard-Jenkins, Tang, Kim, Rivera, Luo, Dong, Straub, Bailey, Eckenhoff, Ma, Pesqueira, Schwesinger, Monge, Yang, Charron, Raina, Parkhi, Borschowa, Moulon, Gupta, Mur-Artal, Pennington, Kulkarni, Miglani, Gondi, Solanki, Diener, Cheng, Green, Saarinen, Patra, Mourikis, Whelan, Singh, Balntas, Baiyya, Dreewes, Pan, Lou, Zhao, Mansour, Zou, Lv, Wang, Yan, Ren, Nardi, and Newcombe], and up to four time-synchronized exocentric views captured via static GoPro cameras. Expert annotators assign one of four discrete proficiency levels—Novice, Early Expert, Intermediate Expert, or Late Expert—serving as ground truth for classification. The label distribution is notably imbalanced, skewed toward intermediate and late experts due to targeted recruitment of skilled participants.
We use all videos with available proficiency estimation labels, following the official benchmark’s training and validation splits. As in prior work [Braun et al.(2025)Braun, Armani, Meier, Moebus, and Holz], we train on the official training set, reserving 10% for validation, and evaluate on the full, held-out official validation set. This setup ensures direct and fair comparability with the accuracy reported in the dataset paper.
The dataset’s rich multi-view structure and diverse activity settings—ranging from kitchens and dance studios to climbing gyms and soccer fields—make it a strong testbed for real-world, multi-view skill assessment.
4.2 Implementation Details
All models were initialized from a TimeSformer backbone [Bertasius et al.(2021)Bertasius, Wang, and Torresani], pre-trained on Kinetics-600 [Carreira and Zisserman(2017)], and fine-tuned for 4 epochs using AdamW (weight decay = 0.01). We used a cosine annealing learning rate schedule with 1 epoch of linear warmup, a fixed batch size of 16, and 16 attention heads. Training was conducted on a single 80GB NVIDIA A100 GPU.
For each video, a fixed number of frames was sampled uniformly from a random temporal window using a frame sampling rate of $n$ (e.g., every $n$-th frame), ensuring temporal coverage within variable-length clips. The selected frames were resized such that the shortest edge is 224 pixels, center-cropped to $224\times 224$, rescaled to the $[0,1]$ range, and normalized with mean $[0.45,0.45,0.45]$ and standard deviation $[0.225,0.225,0.225]$.
Hyperparameters such as the number of input frames, projector hidden size, LoRA rank and scaling, and learning rate were tuned based on the number and type of input views. All other settings were held fixed across experiments (Refer to Section 5.2 and Table 2).
Table 1: Comparison with EgoExo4D proficiency estimation baselines (from [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray]). We report accuracy (%) for egocentric (Ego), exocentric (Exos), and combined views (Ego+Exos). SkillFormer outperforms the baselines in the Exos and Ego+Exos settings with significantly fewer trainable parameters and fewer epochs. Bold denotes the best accuracy; underlined values indicate second-best.
Method Pretrain Ego Exos Ego+Exos Params Epochs
Random - 24.9 24.9 24.9 - -
Majority-class - 31.1 31.1 31.1 - -
TimeSformer - 42.3 40.1 40.8 121M 15
TimeSformer K400 42.9 39.1 38.6 121M 15
TimeSformer HowTo100M 46.8 38.2 39.7 121M 15
TimeSformer EgoVLP 44.4 40.6 39.5 121M 15
TimeSformer EgoVLPv2 45.9 38.0 37.8 121M 15
SkillFormer-Ego K600 45.9 - - 14M 4
SkillFormer-Exos K600 - 46.3 - 20M 4
SkillFormer-EgoExos K600 - - 47.5 27M 4
5 Results
We evaluate SkillFormer on the EgoExo4D [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray] proficiency estimation benchmark using classification accuracy as the primary evaluation metric, in line with the official proficiency demonstrator benchmark protocol. We consider three input configurations: egocentric (Ego, 1 view), exocentric (Exos, 4 views), and multi-view (Ego+Exos, 5 views), enabling us to isolate the contribution of each modality (Sections 5.1 and 5.2). Additionally, we analyze per-domain accuracy to evaluate generalization across diverse real-world tasks (Section 5.3). SkillFormer achieves consistent improvements in both accuracy and efficiency, outperforming multi-view baselines with fewer parameters and significantly lower training cost.
5.1 Overall Accuracy and Efficiency
SkillFormer achieves state-of-the-art classification accuracy in both Exos (46.3%) and Ego+Exos (47.5%) settings, outperforming the best TimeSformer baseline by up to 16.4%. Simultaneously, SkillFormer demonstrates increased computational efficiency by using 4.5x fewer trainable parameters (27M vs. 121M) and requiring 3.75x fewer training epochs (4 vs. 15). These results highlight SkillFormer’s accuracy and compute-efficiency. Random and majority-class baselines perform significantly worse, underscoring the inherent complexity of the task.
It is worth noting that the proficiency label distribution is notably imbalanced, skewed toward intermediate and late experts due to targeted recruitment of skilled participants. This may bias overall accuracy by underrepresenting novice classes.
Table 2: Training hyperparameters and overall accuracy for SkillFormer across view configurations. Fixed across all runs: epochs = 4, batch size = 16, output dim = 768, attention heads = 16. Hid denotes the dimensionality of the intermediate representation used in the CrossViewFusion module. It controls the capacity of the feed-forward transformation and the gating mechanism.
Views Frames LoRA-r LoRA-a Hid LR Params Acc (%)
Ego 32 32 64 1536 5e-5 14M 45.9
Exos 24 48 96 2048 3e-5 20M 46.3
Ego+Exos 16 64 128 2560 2e-5 27M 47.5
5.2 Efficiency and Configuration Analysis
Table 2 details the hyperparameter configurations adopted across different multi-view setups. As the number of views increases, our design prioritizes efficiency without compromising accuracy. To this end, we strategically reduce the number of frames per view—preserving the temporal span with fewer sampled tokens—while proportionally increasing the LoRA rank and hidden dimension in the CrossViewFusion module. This trade-off ensures that the model retains sufficient capacity to capture inter-view dependencies, even under tighter per-view bandwidth.
Our choice of LoRA rank and fusion dimensionality is not arbitrary. Higher ranks allow the adapter to express richer transformations across views, compensating for the reduced token budget. Empirically, we found that increasing these parameters moderately yields significant gains in accuracy while maintaining training efficiency within tractable computational budgets. For example, SkillFormer-Ego+Exos achieves 47.5% accuracy with only 27M trainable parameters—a 4.5x reduction compared to full fine-tuning of the 121M parameter TimeSformer backbone—demonstrating the effectiveness of low-rank adaptation and targeted fusion.
This design reflects a key motivation behind SkillFormer: enabling scalable, parameter-efficient skill recognition across multi-view egocentric and exocentric inputs.
5.3 Per-Domain Accuracy Analysis
Table 3 reports per-scenario accuracy across all models and configurations. SkillFormer consistently outperforms baselines in the Ego+Exos setting for structured and physically grounded activities such as Basketball (77.88%), Cooking (60.53%), and Bouldering (31.87%). These domains benefit from synchronized egocentric and exocentric perspectives, which enable better modeling of spatial layouts and temporally extended actions. The fusion of multi-view signals allows SkillFormer to exploit cross-perspective cues, such as object-hand interactions or full-body movement trajectories, which are often ambiguous or occluded in single-view inputs.
Interestingly, in Music, the Ego-only configuration achieves the highest accuracy (72.41%), suggesting that head-mounted views are sufficient to capture detailed instrument manipulation, while additional views may introduce redundant or misaligned information. This highlights the flexibility of SkillFormer to adapt its fusion mechanism to view-specific signal quality.
Subjective domains like Dancing reveal limitations: SkillFormer’s Ego+Exos accuracy (13.68%) falls significantly below both the majority baseline (51.61%) and the baseline Ego model (55.65%). This indicates that tasks with high intra-class variability and weak structure may not benefit as much from multi-view fusion, or may require additional modalities (e.g., audio) or supervision cues to disambiguate subtle skill indicators.
These trends underscore that SkillFormer is particularly effective in domains requiring precise spatial-temporal reasoning and multi-view integration, aligning with its architectural design focused on efficient yet expressive fusion across diverse camera perspectives.
However, this per-domain evaluation is also affected by class imbalance: domains with few novice samples may exhibit inflated or unstable accuracy due to limited representation of early-stage performers.
Table 3: Per-scenario accuracy (%) for the majority-class baseline, the baseline models from [Grauman et al.(2024)Grauman, Westbury, Torresani, Kitani, Malik, Afouras, Ashutosh, Baiyya, Bansal, Boote, Byrne, Chavis, Chen, Cheng, Chu, Crane, Dasgupta, Dong, Escobar, Forigua, Gebreselasie, Haresh, Huang, Islam, Jain, Khirodkar, Kukreja, Liang, Liu, Majumder, Mao, Martin, Mavroudi, Nagarajan, Ragusa, Ramakrishnan, Seminara, Somayazulu, Song, Su, Xue, Zhang, Zhang, Castillo, Chen, Fu, Furuta, Gonzalez, Gupta, Hu, Huang, Huang, Khoo, Kumar, Kuo, Lakhavani, Liu, Luo, Luo, Meredith, Miller, Oguntola, Pan, Peng, Pramanick, Ramazanova, Ryan, Shan, Somasundaram, Song, Southerland, Tateno, Wang, Wang, Yagi, Yan, Yang, Yu, Zha, Zhao, Zhao, Zhu, Zhuo, Arbelaez, Bertasius, Damen, Engel, Farinella, Furnari, Ghanem, Hoffman, Jawahar, Newcombe, Park, Rehg, Sato, Savva, Shi, Shou, and Wray], and SkillFormer across different view configurations. Bold indicates the best-performing method per scenario, while underlined values represent the second-best. SkillFormer consistently outperforms other models in multi-view setups, particularly in Basketball, Cooking, and Bouldering.
Scenario Majority Baseline SkillFormer
Ego Exos Ego+Exos Ego Exos Ego+Exos
Basketball 36.19 51.43 52.30 55.24 69.03 70.80 77.88
Cooking 50.00 45.00 35.00 35.00 31.58 47.37 60.53
Dancing 51.61 55.65 42.74 42.74 20.51 15.38 13.68
Music 58.97 46.15 69.23 56.41 72.41 68.97 68.10
Bouldering 0.00 25.31 17.28 17.28 30.77 33.52 31.87
Soccer 62.50 56.25 75.00 75.00 70.83 66.67 66.67
6 Limitations and Future Work
Despite strong accuracy and efficiency, SkillFormer has several limitations suggesting future improvements.
Fixed frame sampling may miss critical temporal cues in variable activities, a limitation that could be addressed through adaptive sampling with attention-based keyframe detection. SkillFormer also underperforms significantly in subjective domains like Dancing, suggesting challenges with stylistic variability that domain adaptation strategies such as curriculum learning could improve.
Additionally, the class imbalance in EgoExo4D affects evaluation accuracy, which could be mitigated through class-weighted loss functions or synthetic data generation. Finally, SkillFormer lacks semantic understanding of skill components, limiting interpretability—a challenge that could be addressed through multi-task learning with pose estimation or text-guided supervision. These enhancements would improve SkillFormer’s accuracy in challenging domains while maintaining its computational efficiency.
7 Conclusions
We introduced SkillFormer, a lightweight architecture for multi-view skill assessment that leverages synchronized egocentric and exocentric video inputs. Our approach combines a TimeSformer backbone with a novel CrossViewFusion module, featuring cross-attention, learnable gating, and parameter-efficient LoRA fine-tuning.
SkillFormer makes a dual contribution to multimodal video understanding. First, it achieves state-of-the-art overall classification accuracy on the EgoExo4D proficiency estimation benchmark—particularly in Exos (46.3%) and Ego+Exos (47.5%) settings—outperforming previous baselines by up to 16.4%. Second, it demonstrates strong computational efficiency, using 4.5x fewer trainable parameters and requiring 3.75x fewer training epochs compared to prior approaches.
Our per-domain analysis reveals that SkillFormer excels in structured physical activities that benefit from complementary viewpoints, achieving significantly higher accuracy in Basketball (77.88%), Cooking (60.53%), and Bouldering (33.52%). These improvements are most pronounced when both egocentric and exocentric perspectives provide complementary information about spatial layouts and temporally extended actions.
These results demonstrate that parameter-efficient adaptation and strategic multi-view fusion can substantially enhance skill assessment capabilities while reducing computational demands. SkillFormer’s unified architecture for integrating egocentric and exocentric signals presents a significant advancement in efficient multi-view skill assessment, with particular promise for resource-constrained deployment scenarios.
Acknowledgements
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).